OpenStudy (anonymous):

What is the slope of the line that passes through (3, -2) and (-1, -1)? A.4/1 B.-1/4 C.-4/1 D.1/4

OpenStudy (experimentx):

slope = (y2-y1)/(x2-x1) ... if x1 y1 and x2 y2 are your points.

Directrix (directrix):

[ -1- ( -2) ] / [ -1 - 3] = [-1+2]/ [-4] = 1/[-4] = -1/4